Define the determinant function and write its properties. Also use this definition
Find the determinant of a 3 × 3 matrix.
1
Expert's answer
2020-05-06T16:02:42-0400

The determinant of a function is a function from nxn square matrix to a scalar such that it is defined as:

let a matrix P be defined as P= (aij) where i and j lie between 1 and n.


"\sum_{{1<j<n}}" (-1)i+j a1,j det A1,j where A1,j is a sub-matrix of P.


properties of determinant:

  1. the determinant does not change if rows are changed into columns or columns are changed into rows.( reflection property)
  2. if all the elements of a row or a column are zero, then determinant of the matrix is zero (zero property)
  3. the interchanging of any two rows of a matrix changes the sign.( switching property)
  4. if the row or a column are multiplied by a scale constant then determinant gets multiplied by same scalar constant.( scalar multiple property)
  5. if all the elements of a row/column are same, then determinant is zero.(proportionality)


determinant of 3x3 matrix:


"\begin{vmatrix}
 1 & 2 & 3 \\
 4 & 1 & 2 \\
 2 & 1 & 3 \\
\end{vmatrix}" = 1"\begin{vmatrix}
 1 & 2 \\
 1 & 3
\end{vmatrix}" -2"\begin{vmatrix}
 4 & 2 \\
 2 & 3
\end{vmatrix}" +3"\begin{vmatrix}
 4 & 1 \\
 2 & 1
\end{vmatrix}" = 1-2(8)+3(2) = 1-16+6 = -9